OpenStudy (anonymous):

how would you in numbers explaining get 10^-2 and 11^-2

OpenStudy (agent0smith):

Using exponent rules, a negative exponent can be turned into a positive:\[\huge a^{-b} = \frac{ 1 }{ a^b }\] So...\[\huge 10^{-2} = \frac{ 1 }{ 10^2 }\] Which you can calculate since 10^2 = 10*10 Try the same with 11^-2

OpenStudy (anonymous):

can you show me how

OpenStudy (agent0smith):

\[\huge 11^{-2} = \frac{ 1 }{ 11^2 }\]

OpenStudy (anonymous):

so 1/121

OpenStudy (anonymous):

hello you there

OpenStudy (agent0smith):

1/121 is correct!